OpenStudy (anonymous):

16x^2+128x+81y^2-324y-716=0 Write the equation in the standard form of an ellipse with center (h, k)

OpenStudy (anonymous):

lol oh i see http://www.sosmath.com/algebra/quadraticeq/complsquare/complsquare.html

OpenStudy (anonymous):

Complete the square.

OpenStudy (anonymous):

16x^2+128x+81y^2-324y=716

OpenStudy (anonymous):

ok so were trying to get it in the form (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

you sure thats suppose to be 81y^2?

OpenStudy (anonymous):

Well, it is important to get the exponent of x^2 and y^2 to 1. So first change to this form\[16(x ^{2}+8x)+81(y ^{2}-4)=716\]

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

\[81(y ^{2}-4x)\]

OpenStudy (anonymous):

-4y

OpenStudy (anonymous):

then we get? 16(x^2+8x+16)+81(y^2-4y+4)=736 16(x+4)^2+81(y-2)^2=736

OpenStudy (anonymous):

so from the general form of the equation (x-h)^2+(y-k)^2=r^2 our center (h,k) is (-4,2)

OpenStudy (anonymous):

and you can check it by the general form of a circle which is Ax^2+Ay^2+Bx+Cy+D=0 and the formula for the center is (-b/2a,-c/2a)

OpenStudy (anonymous):

if you have any question feel free to ask but check out the link i first entered thats what helped me

OpenStudy (anonymous):

Wait, wait. Ellipse, we like the RHS to be 1, so divide everything by 736.

OpenStudy (anonymous):

oh okay yes sorry

OpenStudy (anonymous):

Good work, purplec. You're a stud.

OpenStudy (anonymous):

thank you...and stud?..lol

OpenStudy (anonymous):

You're a rockstar!